Bell Media has common stock trading at a price of $74, and a market capitalization of $23 billion. The firm also has preferred stock worth a total of $6 billion, currently trading at $54 per share and paying a dividend of $4.50 per share. The firm's beta is 1.2, the risk-free rate is 2.4%, and the market risk premium is 6%. The firm has $28 billion of debt outstanding. Its bonds with the face value of $10,000 and semi-annual 5% coupons currently have 5 years to maturity and trade at $8783.37. If the firm's tax rate is 30%, what is Bell's WACC?
Answer #1

Cost of Preferred Stock = 4.50/54 = 8.33%

Calculating Cost of Equity using CAPM Model,

Cost of Equity = 0.024 + 1.20(0.06)

Cost of Equity = 9.60%

Calculating Cost of Debt,

Using CAPM Model,

I = [PV = -8,783.37, FV = 10,000, PMT = 250, N = 10]

I = 8.00%

WACC = [23(0.096) + 6(0.0833) + 28(0.08)(1 - 0.30)]/(23 + 6 + 28)

WACC = 7.50%